### v1

# Explicit neural network classifiers for non-separable data

Abstract.
We fully characterize a large class of feedforward neural networks in terms of truncation maps. As an application, we show how a ReLU neural network can implement a feature map which separates concentric data.
1. Introduction
As deep learning models become more powerful and widespread, enhancing their efficiency and interpretability becomes increasingly important. Neural networks, with their billions of parameters trained primarily through gradient descent, often function as black-box models. This opacity makes it challenging to understand how they make decisions. Research on the mathematical foundations of deep learning aims to shed light on these processes, helping to interpret and explain the decision-making mechanisms of neural networks.
In previous work (joint with T. Chen) [CE23, CE24], we showed that a feedforward neural network
$f_{\underline{\theta}}:\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\underline{\theta}_{1}}}{{\longrightarrow}}\,\,\mathbb{R}^{d_{1}}\stackrel{{\scriptstyle\underline{\theta}_{2}}}{{\longrightarrow}}\,\,\cdots\stackrel{{\scriptstyle\underline{\theta}_{L-1}}}{{\longrightarrow}}\mathbb{R}^{d_{L-1}}\stackrel{{\scriptstyle\underline{\theta}_{L}}}{{\longrightarrow}}\mathbb{R}^{d_{L}}$ (1.1)
can be rewritten in terms of endomorphisms $\tau$ of input space, or truncation maps,
$f_{\underline{\theta}}:\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\tau^{(1)}}}{{\longrightarrow}}\,\,\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\tau^{(2)}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\tau^{(L-1)}}}{{\longrightarrow}}\mathbb{R}^{d_{0}}\longrightarrow\mathbb{R}^{d_{L}}.$ (1.2)
This characterization makes it possible to observe how data transforms as it moves along the layers of a neural network. In the case of ReLU activation function, a description of the truncation map in terms of cones yielded a geometric explicit construction of networks which interpolate data for multiclass classification tasks, with greatly reduced number of parameters.
In this work, we continue to develop the truncation map as a tool for investigating the mathematical properties of neural networks. The main technical limitations of previous work were that the width was bounded by the dimension of the input space, and that non-linear decision boundaries for binary classification could not be achieved by the characterization of the truncation map for ReLU. Our contributions are as follows:
  • We extend the description of feedforward neural networks in terms of truncation maps to the case of width $>$ input dimension, and give full characterization of truncation maps for ReLU activation (Section 2).
  • We apply these results to explicitly construct networks which interpolate data in 2 model cases (see Figure 1) for binary classification data which is not linearly separable (Section 3).
The result for concentric data (as in the second picture of Figure 1) is particularly interesting: It realizes the first layer of a ReLU neural network as a piecewise linear version of a feature map $\phi(x)=(x,\left\lVert{x}\right\rVert^{2})$ that a support vector machine might implement. Moreover, it illustrates the fact that ReLU neural networks with width at least (input dimension)$+1$ are universal for real functions [Han19, HS17]. This distribution of data is also one of the examples considered in [Joh19], which proved that functions with a level set containing a bounded component cannot be approximated by ReLU neural networks with width bounded above by the input dimension.
Related work
We make a few remarks on related work beyond what was mentioned in the previous paragraph. First, there are other works with explicit constructions of ReLU neural networks. Notably, [Coo21] gave an explicit construction of a shallow network which interpolates the data. For works characterizing critical points of loss functions for ReLU networks with fixed architecture, we refer the reader to the introduction of [CE24].
Second, we note that the reinterpretation of feedforward ReLU neural networks in terms of truncation maps was independently discovered and studied in [VLL23]. We adopt the nomenclature “polyhedral cones” to match theirs, but otherwise match notation with our previous work [CE23, CE24].
Finally, we observe that the maps $H_{j}$ defined in [HS17] are truncation maps, but the affine maps considered are very restricted, and the focus of that work was not on the geometric structure of those maps.
$\mathcal{X}_{2}$
$\mathcal{X}_{1}$
$v_{1}$
$v_{2}$
$p$
$\mathcal{X}_{1}$
$\mathcal{X}_{2}$
$p$
$v_{1}$
$v_{2}$
$v_{3}$
Figure 1. Simplified representations of the models of data distribution we consider here, and the polyhedral cones constructed such that the truncated data becomes linearly separable. The diagram on the left corresponds to Proposition 3.1, and the diagram on the right corresponds to Corollary 3.2.
2. Setting and technical tools
Consider neural networks of the form
$x^{(0)}$ $=x_{0}\in\mathbb{R}^{d_{0}}\,\,\text{ initial input,}$,
$x^{(\ell)}$ $=\sigma(W_{\ell}x^{(\ell-1)}+b_{\ell})\in\mathbb{R}^{d_{\ell}},\text{ for }\ell=1,\cdots,L-1,$ (2.1)
$x^{(L)}$ $=W_{L}x^{(L-1)}+b_{L}\in\mathbb{R}^{d_{L}},$
where $W_{\ell}\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}$ for $\ell=1,\cdots,L$ are weight matrices, $b_{\ell}\in\mathbb{R}^{d_{\ell}}$ for $\ell=1,\cdots,L$ are bias vectors, and $\sigma$ is an activation function which acts component-wise on vectors. In the present work, we will focus on the ramp function (ReLU),
$\sigma(x)_{i}:=(x_{i})_{+}=\max(x_{i},0).$ (2.2)
Given $W\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$, consider the family of truncation maps
$\tau_{W,b}:\mathbb{R}^{n}$ $\to\mathbb{R}^{n}$
$x$ $\mapsto(W){}^{+}\left(\sigma(Wx+b)-b\right),$ (2.3)
where $(W){}^{+}$ is the generalized inverse. Define the cumulative parameters
$W^{(1)}$ $:=W_{1},\quad W^{(\ell)}$ $:=W_{\ell}\cdots W_{1}=W_{\ell}W^{(\ell-1)},$
$b^{(1)}$ $:=b_{1},\quad\quad b^{(\ell)}$ $:=W_{\ell}b^{(\ell-1)}+b_{\ell},$ (2.4)
associated to a network (2), and for any input $x=x^{(0)}$, let $x^{(\tau,\ell)}$ correspond to the output of $\ell$ many truncation maps applied to $x$,
$x^{(\tau,0)}:=x^{(0)},\quad x^{(\tau,\ell)}$ $:=\tau_{W^{(\ell)},b^{(\ell)}}\,(x^{(\tau,\ell-1)}),\quad\ell=1,\cdots,L.$ (2.5)
In [CE23, CE24], the author and a collaborator showed that a neural network can be described in terms of truncation maps and cumulative parameters: If all weight matrices $W_{\ell}$ are surjective, $\ell=1,\cdots,L$, then
$x^{(\ell)}$ $=W^{(\ell)}x^{(\tau,\,\ell)}+b^{(\ell)},$ (2.6)
for $\ell=1,\cdots,L-1$, and
$x^{(L)}$ $=W^{(L)}x^{(\tau,\,L-1)}+b^{(L)}.$ (2.7)
2.1. Polyhedral cones
In [CE24], the action of a truncation map $\tau$ for a ReLU neural network was given in terms of cones in input space, and $\tau$ acts either as the identity or as a map to a single point on two distinguished regions on input space. This description was used to construct zero loss classifiers for (sequentially) linearly separable data. For data that is not (sequentially) linearly separable, this is not sufficient.
The following lemma completely describes the action of the truncation map with ReLU activation function for surjective weight matrices.
Lemma 2.1.
Given $W:\mathbb{R}^{n}\to\mathbb{R}^{m}$ a surjective matrix and $b\in\mathbb{R}^{m}$, there exist $p\in\mathbb{R}^{n}$ and $(v_{i})_{i=1}^{m}\subset\mathbb{R}^{n}$ such that, for
$x=p+\tilde{x}+\sum_{i=1}^{m}a_{i}v_{i},$ (2.8)
where $\tilde{x}\in\ker W$, then
$\tau_{W,b}(x)$ $=(W){}^{+}Wp+\sum_{i;a_{i}>0}a_{i}v_{i}$
$=(W){}^{+}Wx-\sum_{j;a_{j}\leq 0}a_{j}v_{j}.$ (2.9)
Conversely, given $p\in\mathbb{R}^{n}$ and a linearly independent set $(v_{i})_{i=1}^{m}\subset\mathbb{R}^{n}$, there exist $W\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$ that realize (2.1) for $x\in\mathbb{R}^{n}$ decomposed as
$x=p+\tilde{v}+\sum_{i=1}^{m}a_{i}v_{i},$ (2.10)
where $\tilde{v}\in(span(v_{i})_{i=1}^{m})^{\perp}$.
Proof.
Let $W:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a surjective linear map. Recall that $W(W){}^{+}={\textbf{1}}_{m\times m}$ and $(W){}^{+}W$ is the orthogonal projector to $(\ker W)^{\perp}$. Define
$v_{i}:=(W){}^{+}(e_{i}^{m}),\,\,i=1,\cdots,m,$ (2.11)
and note that $W(v_{i})=W(W){}^{+}e_{i}^{m}=e_{i}^{m}$, and $(W){}^{+}Wv_{i}=(W){}^{+}e_{i}^{m}=v_{i}$. Then fixing $p\in\mathbb{R}^{n}$, for any $x\in\mathbb{R}^{n}$ we have the decomposition
$x=p+\tilde{x}+\sum_{i=1}^{m}a_{i}v_{i},$ (2.12)
where $\tilde{x}\in\ker W$. Let $b=-Wp$. Then
$\tau_{W,b}(x)$ $=(W){}^{+}\left(\sigma(Wx+b)-b\right)$
$=(W){}^{+}\left(\sigma\left(Wp+W\tilde{x}+\sum_{i=1}^{m}a_{i}W(v_{i})-Wp\right)-b\right)$
$=(W){}^{+}\left(\sigma\left(\sum_{i=1}^{m}a_{i}e_{i}\right)+Wp\right)$ (2.13)
$=(W){}^{+}Wp+\sum_{i;a_{i}>0}a_{i}v_{i}$
$=(W){}^{+}Wx-\sum_{j;a_{j}\leq 0}a_{j}v_{j}.$
Note that given $W,b$, where $W$ is surjective, we can find $p:=-(W){}^{+}b$ and $v_{i}:=(W){}^{+}e_{i}^{m}\in\mathbb{R}^{n}$, for $i=1,\cdots,m$. Conversely, given $p\in\mathbb{R}^{n}$ and $(v_{i})_{i=1}^{m}\subset\mathbb{R}^{n}$ linearly independent, we can define $(W){}^{+}:=[v_{1}\cdots v_{n}]$, which is injective and so $W$ is surjective, and $b:=-Wp$. ∎
Consider $W\in GL(n)$ and $b\in\mathbb{R}^{n}$. Then $p\in\mathbb{R}^{n}$ and $\underline{v}:=(v_{1},\cdots,v_{n})$ given by Lemma 2.1 define two polyhedral cones
$\mathcal{S}_{+}(p,\underline{v}):=\left\{p+\sum_{i=1}^{n}a_{i}v_{i}:a_{i}\geq 0,i=1,\cdots,n\right\}$ (2.14)
and
$\mathcal{S}_{-}(p,\underline{v}):=\left\{p+\sum_{i=1}^{n}a_{i}v_{i}:a_{i}\leq 0,i=1,\cdots,n\right\}.$ (2.15)
We will let $\mathcal{S}:=\mathcal{S}_{+}\cup\mathcal{S}_{-}$. Clearly, if $x\in\mathcal{S}_{+}$ then $\tau_{W,b}(x)=x$, and if $x\in\mathcal{S}_{-}$ then $\tau_{W,b}(x)=p.$ Moreover, $\tau_{W,b}(x)\in\mathcal{S}_{+}$ for all $x\in\mathbb{R}^{n}$, and any point in the complement $\mathcal{S}^{c}$ gets projected to the boundary of $\mathcal{S}_{+}$.
$p$
$v_{1}$
$\mathcal{S}_{-}$
$\mathcal{S}_{+}$
$v_{3}$
$v_{2}$
Figure 2. For the ReLU activation function, how a truncation map $\tau$ acts on the data depends on the positioning of the data with respect to the associated polyhedral cone $\mathcal{S}=\mathcal{S}_{+}\cup\,\mathcal{S}_{-}$, characterized by a base point $p$ and edges given by a set of linearly independent vectors $v_{i}$, $i=1,\cdots,m$ (in the picture, $m=3$).
The boundary of $\mathcal{S}_{+}\subset\mathbb{R}^{n}$ can be decomposed into faces, similarly to a simplex, characterized by $a_{k}=0$, $k=1,\cdots,n$. Each one of these $(n-1)$-faces spans a hyperplane in $\mathbb{R}^{n}$,
$H_{k}(p,\underline{v}):=\left\{p+\sum_{i=1,i\neq k}^{n}a_{i}v_{i}:a_{i}\in\mathbb{R}\right\}.$ (2.16)
If $\underline{v}=(v_{1},\cdots,v_{m})$ contains only $m<n$ linearly independent vectors in $\mathbb{R}^{n}$, as may be the case if $W\in\mathbb{R}^{m\times n}$ in Lemma 2.1, then we let
$\mathcal{S}_{+}(p,\underline{v}):=\left\{p+\sum_{i=1}^{m}a_{i}v_{i}+\tilde{v}:a_{i}\geq 0,i=1,\cdots,m,\,\tilde{v}\in(span(v_{i})_{i=1}^{m})^{\perp}\right\},$ (2.17)
and similarly for $\mathcal{S}_{-}$. In this case, $\mathcal{S}(p,\underline{v})$ can be thought of as a family of polyhedral cones $\mathcal{S}(p+\tilde{v},\underline{v})$ defined as in (2.14) on $m$-dimensional affine subspaces
$H_{\tilde{v}}=\left\{\tilde{v}+\sum_{i=1}^{m}a_{i}v_{i}:a_{i}\in\mathbb{R}\right\}\subset\mathbb{R}^{n},$ (2.18)
indexed by $\tilde{v}\in(span(v_{i})_{i=1}^{m})^{\perp}$, or $\tilde{v}\in\ker(W)$ in Lemma 2.1.
We see from (2.1) that two points $x_{1},x_{2}$ for which $\{i,a_{i}(x_{1})>0\}=\{i,a_{i}(x_{2})>0\}=I$ and $a_{i}(x_{1})=a_{i}(x_{2})$ for all such $i\in I$ will be truncated to the same point.
2.2. Increased width
Next, we study what happens if $W$ is not surjective. We leave the case where $W$ is not full rank for future work, and consider here the situation where $W\in\mathbb{R}^{m\times n}$ for $n<m$, and $W$ is injective.
Define
$\iota(n,m)$ $:\mathbb{R}^{n}\to\mathbb{R}^{m},\quad n\leq m,$
$\iota(n,m)(e^{n}_{i})$ $=e^{m}_{i},\quad i=1,\cdots,n,$ (2.19)
where $\left\{e_{i}^{n}\right\}_{i=1}^{n}$ is the standard basis for $\mathbb{R}^{n}$, so that $\iota(n,m)\in\mathbb{R}^{m\times n}$ is a block matrix with upper block ${\textbf{1}}_{n\times n}$ and lower block made of zeros.
Lemma 2.2.
Let $W:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a linear map, for $n<m$. There exists $\tilde{W}\in\mathbb{R}^{m\times m}$ such that $W=\tilde{W}\iota(n,m)$. If $W$ is injective, then $\tilde{W}$ can be made invertible. In that case,
$\sigma(Wx+b)=\tilde{W}\tau_{\tilde{W},b}(\tilde{x})+b,$ (2.20)
for $\tilde{x}=\iota(n,m)x\in\mathbb{R}^{m}$.
Proof.
First, we construct $\tilde{W}$. Suppose $w_{i}\in\mathbb{R}^{m}$, $i=1,\cdots,n$, are the columns of $W$,
$W=\begin{bmatrix}w_{1}&\cdots&w_{n}\end{bmatrix}.$ (2.21)
Then let
$\tilde{W}:=\begin{bmatrix}w_{1}&\cdots&w_{n}&\tilde{w}_{1}&\cdots&\tilde{w}_{m-n}\end{bmatrix},$ (2.22)
so $\tilde{W}(e_{i}^{m})=w_{i}$, for $i=1,\cdots,n$, and $\tilde{W}(e_{n+j}^{m})=\tilde{w}_{j}$ for $j=1,\cdots,m-n$, for an arbitrary collection of vectors $\{\tilde{w}_{j}\}_{j=1}^{m-n}\subset\mathbb{R}^{m}$. We check
$\tilde{W}\iota(n,m)(e_{i}^{n})=\tilde{W}(e_{i}^{m})=w_{i}=W(e_{i}^{n}),$ (2.23)
so $W=\tilde{W}\iota(n,m)$. Moreover, if $W$ is injective, then $\left\{\tilde{w}_{j}\right\}_{j=1}^{m-n}$ can be chosen so that the columns of $\tilde{W}$ are linearly independent, and $\tilde{W}\in GL(n).$
Finally, we check (2.20):
$\tilde{W}\tau_{\tilde{W},b}(\iota(n,m)x)+b$ $=\tilde{W}(\tilde{W}){}^{+}\left(\sigma(\tilde{W}\iota(n,m)x+b)-b\right)+b$
$=\sigma\left(Wx+b\right).$ (2.24)
Thus, an increase in width can be split into a linear layer which includes into a higher-dimensional space $\iota:\mathbb{R}^{n}\hookrightarrow\mathbb{R}^{m}$, followed by a non-linear map between layers with the same dimension.
We can now rewrite a neural network for which $d_{0}<d_{1}$ and $d_{1}\geq d_{2}\geq\cdots\geq d_{L}$ in terms of truncation maps: Consider such a neural network
$\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\underline{\theta}_{1}}}{{\longrightarrow}}\,\,\mathbb{R}^{d_{1}}\stackrel{{\scriptstyle\underline{\theta}_{2}}}{{\longrightarrow}}\,\,\mathbb{R}^{d_{2}}\stackrel{{\scriptstyle\underline{\theta}_{3}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\underline{\theta}_{L}}}{{\longrightarrow}}\mathbb{R}^{d_{L}}$ (2.25)
with $L$ layers and parameters $\underline{\theta}=(\underline{\theta}_{1},\cdots,\underline{\theta}_{L}$), with $\underline{\theta}_{\ell}=(W_{\ell},b_{\ell})\in\mathbb{R}^{d_{\ell}\times d_{\ell-1}}\times\mathbb{R}^{d_{\ell}}$ for $\ell=1,\cdots,L$. Assuming all weight matrices are full-rank, and letting $\tilde{x}=\iota(d_{0},d_{1})(x)$, (2.20) gives
$x^{(1)}$ $=\sigma(W_{1}x+b_{1})=\sigma(\tilde{W}_{1}(\tilde{x})+b_{1})=\tilde{W}_{1}\tau_{\tilde{W}_{1},b_{1}}(\tilde{x})+b_{1}$ (2.26)
for a suitable $\tilde{W}_{1}\in GL(d_{1})$, which implies
$x^{(2)}$ $=\sigma(W_{2}x^{(1)}+b_{2})$
$=\sigma(W_{2}\tilde{W}_{1}\tau_{\tilde{W}_{1},b_{1}}(\tilde{x})+W_{2}b_{1}+b_{2})$
$=\sigma(\tilde{W}^{(2)}\tilde{x}^{(\tau,1)}+b^{(2)})$ (2.27)
$=\tilde{W}^{(2)}\tilde{x}^{(\tau,2)}+b^{(2)},$
where we took $\tilde{W}^{(2)}=W_{2}\tilde{W}_{1}$ instead of $W^{(2)}=W_{2}W_{1}$ as in the $d_{0}\geq d_{1}$ case. Then a modified version of (2.7) holds,
$x^{(L)}=\tilde{W}^{(L)}\tilde{x}^{(\tau,L-1)}+b^{(L)},$ (2.28)
where
$\tilde{x}^{(\tau,L-1)}=\tau_{\tilde{W}^{(L-1)},b^{(L-1)}}\circ\cdots\circ\tau_{\tilde{W}^{(1)},b^{(1)}}(\,\tilde{x}\,),$ (2.29)
for
$\tilde{W}^{(\ell)}=W_{\ell}\cdots W_{2}\tilde{W}_{1},\quad\ell=1,\cdots,L.$ (2.30)
Finally, (2.28) implies that (2.25) is equivalent to
$\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\iota}}{{\hookrightarrow}}\mathbb{R}^{d_{1}}\stackrel{{\scriptstyle\tau_{\underline{\theta}^{(1)}}}}{{\longrightarrow}}\,\,$ $\mathbb{R}^{d_{1}}\stackrel{{\scriptstyle\tau_{\underline{\theta}^{(2)}}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\tau_{\underline{\theta}^{(L-1)}}}}{{\longrightarrow}}\mathbb{R}^{d_{1}}\stackrel{{\scriptstyle\underline{\theta}^{(L)}}}{{\longrightarrow}}\mathbb{R}^{d_{L}},$ (2.31)
where $\underline{\theta}^{(\ell)}=(\tilde{W}^{(\ell)},b^{(\ell)})$, for $\ell=1,\cdots,L$.
Remark 2.3.
While Lemma 2.2 could be applied at any given layer of a neural network, and (with extra work) could be used multiple times so that a version of (2.28) would hold for arbitrary feedforward networks, we restrict ourselves to the case $d_{0}\leq d_{1}\geq d_{2}\geq\cdots\geq d_{L}$ in this paper because it is natural, and sufficient for dealing with the toy models considered here. Moreover, it encompasses the architectures needed to approximate continuous functions on a compact set [HS17].
3. Application to classification problems
We will be considering classification tasks, so that a dataset $\mathcal{X}=\sum_{j=1}^{Q}\mathcal{X}_{j}$ is split into $Q$ disjoint classes. For $\mathcal{X}\subset\mathbb{R}^{d_{0}}$, a classifying neural network is a map
$f_{\underline{\theta}}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{Q}$ (3.1)
parametrized by a vector of parameters $\underline{\theta}$.
In the following results, we show how one truncation map can make the data (sequentially) linearly separable for 3 toy models of data that is not linearly separable, with $Q=2$ and $3$. Then the results from [CE24] can be applied to construct a neural network which interpolates the data.
We start with the case where the data can be separated by a single polyhedral cone, as in the first image in Figure 1.
Proposition 3.1.
Suppose $\mathcal{X}=\mathcal{X}_{1}\cup\mathcal{X}_{2}\subset\mathbb{R}^{n}$, and there exist $p\in\mathbb{R}^{n}$ and $\{v_{i}\}_{i=1}^{m}\subset\mathbb{R}^{n}$ a set of linearly independent vectors, such that $\mathcal{X}_{1}\subset\mathcal{S}_{-}(p,\underline{v})$ and $\mathcal{X}_{2}\subset(\mathcal{S}_{-})^{c}$, for $n\geq m$, and $\underline{v}=(v_{i})_{i=1}^{m}$.
Then there exist $W\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$ such that the truncated data $\tau_{W,b}(\mathcal{X})$ is linearly separable.
Proof.
Let $W$ and $b$ be defined as in Lemma 2.1. For all $x\in\mathcal{X}_{1}$ we have
$\tau_{W,b}(x)=(W){}^{+}Wp.$ (3.2)
For any $x\in\mathcal{X}_{2}$, written as
$x=p+\tilde{x}+\sum_{i=1}^{m}a_{i}(x)v_{i},$ (3.3)
for some $\tilde{x}\in\ker W$, there exists $i\in\left\{1,\cdots,n\right\}$ such that $a_{i}(x)>0$, and
$\tau_{W,b}(x)=(W){}^{+}Wp+\sum_{i;a_{i}(x)>0}a_{i}(x)\,v_{i}.$ (3.4)
We can now construct a hyperplane which separates the two classes. Let
$\tilde{a}:=\min\left\{a_{i}(x):x\in\mathcal{X}_{2},i=1,\cdots,m,\text{ and }a_{i}(x)>0\right\},$ (3.5)
and consider the $(m-1)$-dimensional affine subspace
$H(p,(v_{i})_{i=1}^{m},\tilde{a}):=\left\{(W){}^{+}Wp+\frac{\tilde{a}}{2}v_{1}+\sum_{j=2}^{m}h_{j}(v_{1}-v_{j}):h_{j}\in\mathbb{R},j=2,\cdots,m\right\}\subset\mathbb{R}^{n},$ (3.6)
which we will denote simply by $H$ in the remainder of this proof. Note that $H,\tau_{W,b}(\mathcal{X}_{1})$ and $\tau_{W,b}(\mathcal{X}_{2})\subset(\ker W)^{\perp}$. Then for $x\in\mathcal{X}_{1}$,
$\tau_{W,b}(x)$ $=(W){}^{+}Wp$
$=\left((W){}^{+}Wp+\frac{\tilde{a}}{2}v_{1}\right)-\frac{\tilde{a}}{2}v_{1},$ (3.7)
and for $x\in\mathcal{X}_{2}$,
$\tau_{W,b}(x)$ $=(W){}^{+}Wp+\sum_{a_{i}(x)>0}a_{i}(x)v_{i}$
$=(W){}^{+}Wp+\sum_{i=1}^{m}\sigma(a_{i}(x))v_{i}$
$=(W){}^{+}Wp+\frac{\tilde{a}}{2}v_{1}+\left(\sigma(a_{1}(x))-\frac{\tilde{a}}{2}\right)v_{1}+\sum_{i>1}\sigma(a_{i}(x))v_{i}$ (3.8)
$=\left((W){}^{+}Wp+\frac{\tilde{a}}{2}v_{1}+\sum_{i>1}\sigma(a_{i}(x))(v_{i}-v_{1})\right)+\underbrace{\left(\sum_{i=1}^{m}\sigma(a_{i}(x))-\frac{\tilde{a}}{2}\right)}_{>0}v_{1},$
from which we see that $\tau_{W,b}(\mathcal{X}_{1})$ and $\tau_{W,b}(\mathcal{X}_{2})$ are on opposite sides of the hyperplane $H$ inside $(\ker W)^{\perp}$. To extend $H$ to a hyperplane in $\mathbb{R}^{n}$ which separates the truncated classes, it suffices to take $H+\ker W$. ∎
In this case, the data can be interpolated by applying a second truncation map and an affine map,
$\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\tau_{\underline{\theta}^{(1)}}}}{{\longrightarrow}}\,\,\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\tau_{\underline{\theta}^{(2)}}}}{{\longrightarrow}}\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\underline{\theta}^{(3)}}}{{\longrightarrow}}\mathbb{R}^{2},$ (3.9)
which is equivalent to a neural network with the following architecture:
$\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\underline{\theta}_{1}}}{{\longrightarrow}}\mathbb{R}^{d_{1}}\stackrel{{\scriptstyle\underline{\theta}_{2}}}{{\longrightarrow}}\mathbb{R}^{d_{2}}\stackrel{{\scriptstyle\underline{\theta}_{3}}}{{\longrightarrow}}\mathbb{R}^{2},$ (3.10)
where $d_{0}\geq d_{1}\geq d_{2}\geq 2$.
Consider now data which is concentric, as in the second image of Figure 1. It is easy to see in the 2-dimensional case that no polyhedral cone in $\mathbb{R}^{2}$ will satisfy the assumptions of Proposition 3.1. By adding one extra dimension and using a cone in $\mathbb{R}^{3}$ (as in the middle image of Figure 1), the data can be projected in a way that makes it linearly separable.
Corollary 3.2.
Suppose $\mathcal{X}=\mathcal{X}_{1}\cup\mathcal{X}_{2}\subset\mathbb{R}^{n}$, and there exists an $n$-simplex $S$ such that $\mathcal{X}_{1}\subset S$ and $\mathcal{X}_{2}\subset S^{c}$. Then there exist $W\in GL(n+1)$ and $b\in\mathbb{R}^{n+1}$ such that $\tau_{W,b}\circ\iota(n,n+1)(\mathcal{X})$ is linearly separable.
Proof.
First, include the data into $\mathbb{R}^{n+1}$ with $\iota:=\iota(n,n+1)$. Next, let $s_{1},\cdots,s_{n+1}$ be the vertices for $S^{\prime}=\iota(S)$, and choose $p\in\mathbb{R}^{n+1}$ a positive distance away from the hyperplane in $\mathbb{R}^{n+1}$ spanned by $S^{\prime}$. Let
$v_{i}:=p-s_{i},\quad i=1,\cdots,n+1.$ (3.11)
Then a polyhedral cone $\mathcal{S}_{-}$ can be constructed from $p$ and $\left(v_{i}\right)_{i=1}^{n+1}$ for which the assumptions of Proposition 3.1 are satisfied: The simplex $S^{\prime}$ is entirely contained in $\mathcal{S}_{-}$ by construction, so that
$\iota(\mathcal{X}_{1})\subset S^{\prime}\subset\mathcal{S}_{-}.$ (3.12)
Moreover, $\iota(\mathcal{X}_{2})\subset(\mathcal{S}_{-})^{c}$, since $\iota(\mathcal{X}_{2})\subset(S^{\prime})^{c}\cap span(S^{\prime})$ and $\mathcal{S}_{-}\cap span(S^{\prime})=S^{\prime}$.
Proposition 3.1 now applies, and a hyperplane like (3.6) can be found that separates $\tau_{W,b}(\iota(\mathcal{X}_{1}))$ and $\tau_{W,b}(\iota(\mathcal{X}_{2}))$, for $W\in GL(n+1)$ and $b\in\mathbb{R}^{n+1}$ given by Lemma 2.1. ∎
Putting together Lemma 2.2 and Corollary 3.2 gives our main result:
Theorem 3.3.
Suppose $\mathcal{X}=\mathcal{X}_{1}\cup\mathcal{X}_{2}\subset\mathbb{R}^{d_{0}}$, and there exists a $d_{0}$-simplex $S$ such that $\mathcal{X}_{1}\subset S$ and $\mathcal{X}_{2}\subset S^{c}$. Moreover, let the classes be labeled by two linearly independent vectors $y_{1},y_{2}\in\mathbb{R}^{2}$. Then this data can be interpolated by a feedforward ReLU neural network (2) with $2$ hidden layers,
$\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\underline{\theta}_{1}}}{{\longrightarrow}}\mathbb{R}^{d_{1}}\stackrel{{\scriptstyle\underline{\theta}_{2}}}{{\longrightarrow}}\mathbb{R}^{d_{2}}\stackrel{{\scriptstyle\underline{\theta}_{3}}}{{\longrightarrow}}\mathbb{R}^{2},$ (3.13)
where $d_{1}=d_{0}+1$ and $d_{1}\geq d_{2}\geq 2$.
Proof.
We will construct the following sequence of maps to interpolate the data:
$\mathbb{R}^{d_{0}}\stackrel{{\scriptstyle\iota}}{{\hookrightarrow}}\mathbb{R}^{d_{0}+1}\stackrel{{\scriptstyle\tau_{\underline{\theta}^{(1)}}}}{{\longrightarrow}}\,\,\mathbb{R}^{d_{0}+1}\stackrel{{\scriptstyle\tau_{\underline{\theta}^{(2)}}}}{{\longrightarrow}}\mathbb{R}^{d_{0}+1}\stackrel{{\scriptstyle\underline{\theta}^{(3)}}}{{\longrightarrow}}\mathbb{R}^{2}.$ (3.14)
Corollary 3.2 gives the inclusion map $\iota:=\iota(d_{0},d_{0}+1)$, and the parameters $\tilde{W}_{1}\in GL(d_{0}+1)$ and $b^{(1)}\in\mathbb{R}^{d_{0}+1}$ for $\underline{\theta}^{(1)}=(\tilde{W}_{1},b^{(1)})$, such that $\tau_{\underline{\theta}^{(1)}}(\iota(\mathcal{X}))$ is linearly separable. Then the cumulative parameters $\underline{\theta}^{(2)}=(W^{(2)},\,b^{(2)})$ defining the second truncation map and $\underline{\theta}^{(3)}=(W^{(3)},\,b^{(3)})$ defining the affine map in (3.14) can be found using the hyperplane given by (3.6) for Corollary 3.2 and the construction in [CE24].
Finally, the parameters $(\underline{\theta}_{i})_{i=1}^{3}$ defining (3.13) are given according to Lemma 2.2 and (2.30): The weight matrices are
$W_{1}$ $=\tilde{W}_{1}\iota(d_{0},d_{0}+1),$
$W_{2}$ $=W^{(2)}\tilde{W}_{1}^{-1},$ (3.15)
$W_{3}$ $=W^{(3)}(W^{(2)}){}^{+},$
the bias vectors are
$b_{1}$ $=b^{(1)},$
$b_{2}$ $=b^{(2)}-W_{2}b^{(1)},$ (3.16)
$b_{3}$ $=b^{(3)}-W_{3}b^{(2)},$
and $\underline{\theta}_{i}=(W_{i},b_{i})$, $i=1,2,3$. ∎
Acknowledgments: The author thanks Thomas Chen for helpful comments, and J. Elisenda Grigsby for helpful discussions and for pointing out the the works by Hanin and Sellke, and Johnson. P.M.E. was supported by NSF grant DMS-2009800.